# Embedded structure in quantum theory, functional operator and multiverse

### v2

@@ 4 Universes with different particle contents
Now it is time to tackle Q3, i.e., “Why is our universe described by the standard model at the electroweak scale?”. Under the precondition that the existence of our universe can be understood by the combination of the level II multiverse (an ensemble of foreign universes) and the anthropic principle and the above question is not really acknowledged as a problem, we face the question whether a framework to describe the level II multiverse can be constructed or not. In the following, we investigate a theoretical framework to deal with a set of universes with different elementary particles and parameters based on the physical laws of QFT.
@@ 4.1 Level II multiverse
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\widehat{\varPhi}_{k^{(a_{1})}}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{k^{(a_{M})}}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle,$ (94)
where the summation is taken in the range $a_{1}>\cdots>a_{M}$,555 When a label $a_{n}$ takes continuous values, a summation should be replaced into an integration. Same applies to the following. $t$ is an auxiliary parameter which is identified as a time when a system is limited in some observable region of our universe, ${\varPsi}\left(\{\varphi_{k^{(a_{1})}}\},\cdots,\{\varphi_{k^{(a_{M})}}\},t\right)$ is a state’s functional related to $M$ kinds of universes, and $\widehat{\varPhi}_{k^{(a_{1})}}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{k^{(a_{M})}}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle$ is a vacuum state described by
$\widehat{\varPhi}_{k^{(a_{1})}}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{k^{(a_{M})}}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle=|0\rangle_{\{\varphi_{k^{(a_{1})}}\}}\otimes\cdots\otimes|0\rangle_{\{\varphi_{k^{(a_{M})}}\}}.$ (95)
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{1}{\sqrt{M_{(a_{1})}!}}\left(\widehat{\varPhi}_{k^{(a_{1})}}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\right)^{M_{(a_{1})}}\cdots\frac{1}{\sqrt{M_{(a_{\mathscr{N}})}!}}\left(\widehat{\varPhi}_{k^{(a_{\mathscr{N}})}}^{\dagger}\left(\{\varphi_{k^{(a_{\mathscr{N}})}}\}\right)\right)^{M_{(a_{\mathscr{N}})}}|0\!\!\!\!\!\big{/}\rangle.$ (96)
@@
In the Heisenberg picture, installation and removal operators are given in a form with time-dependence by
@@ 4.2.1 Inflation
$\widehat{\varPhi}_{(a)}\left(\{\varphi_{k^{(a)}}\},t\right)=\widehat{\varPhi}_{(a)}\left(\{\varphi_{k^{(a)}}\}\right)e^{-\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t},~{}~{}\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\},t\right)=\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\}\right)e^{\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t},$ (107)
where $\mathscr{E}_{0}^{(a)}=\langle 0|\widehat{H}_{\{\varphi_{k^{(a)}}\}}|0\rangle$ is the vacuum energy of the universe labeled by $a$. Then, the vacuum state is evolved as
$\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\},t\right)|0\!\!\!\!\!\big{/}\rangle=\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\}\right)e^{\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t}|0\!\!\!\!\!\big{/}\rangle=e^{\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t}|0\rangle_{\{\varphi_{k^{(a)}}\}}.$ (108)
Here, we consider a universe with a uniformity and isotropy, whose geometry and dynamics are described by the Robertson-Walker metric. When the vacuum energy density $\rho_{\rm V}^{(a)}\equiv\mathscr{E}_{0}^{(a)}/V$ ($V$:the volume of the universe) dominates over other energy densities, the scale factor $a(t)$ varies based on the Friedmann equation:
$\left(\frac{\dot{a}}{a}\right)^{2}=\frac{8\pi G_{\rm N}}{3}\rho_{\rm V}^{(a)},$ (109)
where $\dot{a}\equiv{da(t)}/{dt}$ and $G_{\rm N}$ is gravitational constant. Then, in the case with $\rho_{\rm V}^{(a)}>0$, an inflation (an exponential expansion of the universe) [11] occurs as
@@ 4.2.2 Third quantization
In a universe whose spacetime varies, the spacetime itself is regarded as a dynamical object, and a theory must be invariant under the general transformation of coordinates and contain a gravitational field.
@@ 4.2.3 Landscape
As a mechanism to produce the level II multiverse, the eternal inflation using the potential $\widehat{V}_{\{\varphi_{\rm L}\}}$ in the landscape looks promising. Then, some of $\varphi_{{\rm L}m}$ can play a role of inflatons. For more detail, a vacuum state $|0\rangle_{\{\varphi_{\rm L}\}}$ and a spacetime can emerge by the operation of $\widehat{\varPhi}_{\rm L}\left(\{\varphi_{\rm L}\},t\right)$ such as $|0\rangle_{\{\varphi_{\rm L}\}}=\widehat{\varPhi}_{\rm L}\left(\{\varphi_{\rm L}\},t\right)|0\!\!\!\!\!\big{/}\rangle$. Each region of the space expands while $\varphi_{{\rm L}m}$ are rolling down towards a local minimum of the potential. When the potential has flat directions for some $\varphi_{{\rm L}m}$, those values can correspond to values of parameters. After inflatons arrive at the local minimum labeled by $\{\varphi_{\rm L}^{(a)}\}$ including a set of parameters, the values of $\varphi_{{\rm L}m}$ are fixed, the vacuum $|0\rangle_{\{\varphi_{\rm L}\}}$ changes into $|0\rangle_{\{\varphi_{k^{(a)}}\}}$, and a universe with $|0\rangle_{\{\varphi_{k^{(a)}}\}}$ comes into existence. If the universe has a positive vacuum energy coming from an inflaton in $\{\varphi_{k^{(a)}}\}$, an inflation can occurs. After the inflation comes to an end, the inflaton decays into other particles in $\{\varphi_{k^{(a)}}\}$ and the universe becomes extremely hot and dense (like a big bang in our universe). In this way, various universes with different particle contents and parameters (different physical laws) can be generated.
A change between universes with different physical laws can also occur in the setup based on the action integral eq. (118), if there are no selection rules to forbid the change. According to the laws of quantum physics, we can calculate the transition amplitude $\widehat{S}_{\rm f~{}\!i}$ from a vacuum state in a universe with $\{\varphi_{k^{(a)}}\}$ to a different one with $\{\varphi_{k^{(b)}}\}$ by using the formula:
@@ 5 Conclusions and discussions
Taking the embedded structure in QFT as a guiding principle, we have arrived at a wider theoretical framework called quantum field’s functional (QFFT) that has QFT built-in. QFFT is constructed by using functional operators and has the nested structure666 Another extension of QFT owning a nested structure has been carried out by promoting a wave functional in QFT into an operator in reference to the second quantization [22]. In this formalism, the operator can create or annihilate a multi-particle state..
@@
On the one hand, QFFT would not be well-defined mathematically in the same way as QFT, that is, it is not yet known whether functional operators can be rigorously defined or not. On the other hand, QFFT has several advantages. First, it is relatively naturally understood that fermions (fields as Grassmann variables) and bosons are represented by field operators satisfying the anti-commutation relations and the commutation relations, respectively. Second, the which-came-first-particles-or-spacetime problem (which-came-first-physical-laws-or-our-universe problem) can be solved, if particles and space-time (physical laws and our universe) can be installed as a unit by a functional operator. Third, the level II multiverse could be described by QFFT. Last, topics related to a beginning of the universe such as inflation, the third quantization and the landscape can be studied in our formulation. Our framework might be also applied to the level III multiverse where the level III multiverse means quantum mechanical many worlds [21]777 It is proposed that the eternally inflating multiverse and quantum mechanical many worlds are the same thing [23]..
@@ Appendix B Wave functional
$~{}~{}~{}+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})\langle\varphi|0\rangle.$ (139)
@@
$e^{-\frac{i}{\hbar}\mathscr{E}_{N}t}C_{N}(\varphi)=\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\varphi^{\dagger}(\bm{x}_{N})\cdots\varphi^{\dagger}(\bm{x}_{N}).$ (141)
@@
$~{}\int\mathscr{D}\varphi~{}\varphi(\bm{x}_{1})\cdots\varphi(\bm{x}_{N})\varphi^{\dagger}(\bm{x}^{\prime}_{1})\cdots\varphi^{\dagger}(\bm{x}^{\prime}_{N})\mathcal{V}_{0}^{\dagger}(\varphi)\mathcal{V}_{0}(\varphi)$
@@
$~{}\mathcal{V}_{0}(\varphi)\mathcal{V}^{\dagger}_{0}(\widetilde{\varphi})+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})\widetilde{\varphi}(\bm{x}_{1})\cdots\widetilde{\varphi}(\bm{x}_{N})\mathcal{V}_{0}(\varphi)\mathcal{V}^{\dagger}_{0}(\widetilde{\varphi})$
@@
$\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)=\int\mathscr{D}\varphi~{}~{}\varphi(\bm{x}_{1})\cdots\varphi(\bm{x}_{N})~{}\mathcal{V}_{0}^{\dagger}(\varphi)\varPsi(\varphi,t).$ (147)
@@
$|\varPsi(t)\rangle$ $=\psi^{(0)}(t)|0\rangle+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\widehat{\varphi}^{\dagger}(\bm{x}_{1})\cdots\widehat{\varphi}^{\dagger}(\bm{x}_{N})|0\rangle$
@@
$~{}+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\widehat{\varphi}^{\dagger}(\bm{x}_{1})\cdots\widehat{\varphi}^{\dagger}(\bm{x}_{N})\int\mathscr{D}\varphi|\varphi\rangle\langle\varphi|0\rangle$
$=\psi^{(0)}(t)|0\rangle+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})|0\rangle$
@@
${\varPsi}(\{\varphi\},t)\equiv\psi^{(0)}(t)+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N}).$ (149)

### v3

@@ 4 Universes with different particle contents
Now it is time to tackle Q3, i.e., “Why is our universe described by the standard model at the electroweak scale?”. Under the precondition that the existence of our universe can be understood by the combination of the level II multiverse (an ensemble of foreign universes) and the anthropic principle, and the above question is not really acknowledged as a problem, we face the question whether a framework to describe the level II multiverse can be constructed or not. In the following, we investigate a theoretical framework to deal with a set of universes with different elementary particles and parameters based on the physical laws of QFT.
@@ 4.1 Level II multiverse
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\widehat{\varPhi}_{(a_{1})}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{(a_{M})}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle,$ (103)
where the summation is taken in the range $a_{1}>\cdots>a_{M}$,555 When a label $a_{n}$ takes continuous values, a summation should be replaced into an integration. Same applies to the following. $t$ is an auxiliary parameter which is identified as a time when a system is limited in some observable region of our universe, ${\varPsi}\left(\{\varphi_{k^{(a_{1})}}\},\cdots,\{\varphi_{k^{(a_{M})}}\},t\right)$ is a state’s functional related to $M$ kinds of universes, and $\widehat{\varPhi}_{(a_{1})}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{(a_{M})}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle$ is a vacuum state described by
$\widehat{\varPhi}_{(a_{1})}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\cdots\widehat{\varPhi}_{(a_{M})}^{\dagger}\left(\{\varphi_{k^{(a_{M})}}\}\right)|0\!\!\!\!\!\big{/}\rangle=|0\rangle_{\{\varphi_{k^{(a_{1})}}\}}\otimes\cdots\otimes|0\rangle_{\{\varphi_{k^{(a_{M})}}\}}.$ (104)
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{1}{\sqrt{M_{(a_{1})}!}}\left(\widehat{\varPhi}_{(a_{1})}^{\dagger}\left(\{\varphi_{k^{(a_{1})}}\}\right)\right)^{M_{(a_{1})}}\cdots\frac{1}{\sqrt{M_{(a_{\mathscr{N}})}!}}\left(\widehat{\varPhi}_{(a_{\mathscr{N}})}^{\dagger}\left(\{\varphi_{k^{(a_{\mathscr{N}})}}\}\right)\right)^{M_{(a_{\mathscr{N}})}}|0\!\!\!\!\!\big{/}\rangle.$ (105)
@@
In the Heisenberg picture, installation and removal operators are given in a form with the time-dependence by
@@ 4.2.1 Inflation
$\widehat{\varPhi}_{(a)}\left(\{\varphi_{k^{(a)}}\},t\right)=e^{-\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t}\widehat{\varPhi}_{(a)}\left(\{\varphi_{k^{(a)}}\}\right),~{}~{}\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\},t\right)=\widehat{\varPhi}^{\dagger}_{(a)}\left(\{\varphi_{k^{(a)}}\}\right)e^{\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t},$ (116)
where $\mathscr{E}_{0}^{(a)}=\langle 0|\widehat{H}_{\{\varphi_{k^{(a)}}\}}|0\rangle$ is the vacuum energy of the universe labeled by $a$. And the vacuum wave function $\varPsi_{\rm V}^{(a)}(t)$ is evolved as
$\varPsi_{\rm V}^{(a)}(t)=e^{-\frac{i}{\hbar}\mathscr{E}_{0}^{(a)}t}\varPsi_{\rm V}^{(a)}(0).$ (117)
Here, we consider a universe labeled by $a$ with a uniformity and isotropy, whose geometry and dynamics are described by the Robertson-Walker metric. When the vacuum energy density $\rho_{\rm V}^{(a)}\equiv\mathscr{E}_{0}^{(a)}/V$ ($V$:the volume of the universe) dominates over other energy densities, the scale factor $a(t)$ varies based on the Friedmann equation:
$\left(\frac{\dot{a}(t)}{a(t)}\right)^{2}=\frac{8\pi G_{\rm N}}{3}\rho_{\rm V}^{(a)},$ (118)
where $\dot{a}(t)\equiv{da(t)}/{dt}$ and $G_{\rm N}$ is gravitational constant. Then, in the case with $\rho_{\rm V}^{(a)}>0$, an inflation (an exponential expansion of the universe) [11] occurs as
@@ 4.2.2 Third quantization
In a universe labeled by $a$ whose spacetime varies, the spacetime itself is regarded as a dynamical object, and a theory must be invariant under the general transformation of coordinates and contain a gravitational field.
@@ 4.2.3 Landscape
As a mechanism to produce the level II multiverse, the eternal inflation using the potential $\widehat{V}_{\{\varphi_{\rm L}\}}$ in the landscape looks promising. Then, some of $\varphi_{{\rm L}m}$ can play a role of inflatons. For more detail, a vacuum state $|0\rangle_{\{\varphi_{\rm L}\}}$ and a spacetime can emerge by the operation of $\widehat{\varPhi}^{\dagger}_{\rm L}\left(\{\varphi_{\rm L}\},t\right)$ such as $|0\rangle_{\{\varphi_{\rm L}\}}=\widehat{\varPhi}^{\dagger}_{\rm L}\left(\{\varphi_{\rm L}\},t\right)|0\!\!\!\!\!\big{/}\rangle$. Each region of the space expands while $\varphi_{{\rm L}m}$ are rolling down towards a different local minimum of the potential or they are transitioning from a local minimum to another one. When the potential has flat directions for some $\varphi_{{\rm L}m}$, those values can correspond to values of parameters. After $\varphi_{{\rm L}m}$ arrive at the local minimum labeled by $\{\varphi_{\rm L}^{(a)}\}$ including a set of parameters, the values of $\varphi_{{\rm L}m}$ are fixed, the vacuum $|0\rangle_{\{\varphi_{\rm L}\}}$ changes into $|0\rangle_{\{\varphi_{k^{(a)}}\}}$, and a universe with $|0\rangle_{\{\varphi_{k^{(a)}}\}}$ comes into existence. If the universe is dominated by a positive vacuum energy coming from some particle (called an inflaton) in $\{\varphi_{k^{(a)}}\}$, an inflation can occurs in the universe. After the inflation comes to an end, the inflaton decays into other particles in $\{\varphi_{k^{(a)}}\}$ and the universe becomes extremely hot and dense (like a big bang in our universe). In this way, various universes with different particle contents and parameters (different physical laws) can be generated.
A change between universes with different physical laws can also occur and its rate can be estimated in the setup based on the action integral (127), if there are no selection rules to forbid the change. According to the laws of quantum physics, we can calculate the transition amplitude $\widehat{S}_{\rm f~{}\!i}$ from a vacuum state in a universe with $\{\varphi_{k^{(a)}}\}$ to a different one with $\{\varphi_{k^{(b)}}\}$ by using the formula:
@@ 5 Conclusions and discussions
Taking the embedded structure in QFT as a guiding principle, we have arrived at a wider theoretical framework called quantum field’s functional (QFFT) that has QFT built-in. QFFT is constructed by using functional operators and has the nested structure666 Another extension of QFT owning a nested structure has been carried out by promoting a wave functional in QFT into an operator in reference to the second quantization [22]. In this formulation, the operator can create or annihilate a multi-particle state..
@@
On the one hand, QFFT would not be well-defined mathematically in the same way as QFT, that is, it is not yet known whether functional operators can be rigorously defined or not. On the other hand, QFFT has several advantages. First, it is relatively naturally understood that fermions (fields as Grassmann variables) and bosons are represented by field operators satisfying the anti-commutation relations and the commutation relations, respectively. Second, the which-came-first-particles-or-spacetime problem (which-came-first-physical-laws-or-our-universe problem) can be solved, if particles and space-time (physical laws and our universe) can be installed as a unit by a functional operator. Third, the level II multiverse could be described by QFFT. Last, topics related to a beginning of the universe such as an inflation, the third quantization and the landscape can be studied in our formulation. Our framework might also be applied to the level III multiverse where the level III multiverse means quantum mechanical many worlds [21]777 It is proposed that the eternally inflating multiverse and quantum mechanical many worlds are the same thing [23]..
@@ Appendix B Wave functional
$~{}~{}~{}+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})\langle\varphi|0\rangle.$ (148)
@@
$e^{-\frac{i}{\hbar}\mathscr{E}_{N}t}C_{N}(\varphi)=\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N}).$ (150)
@@
$~{}\int\mathscr{D}\varphi~{}\frac{1}{N!}\varphi(\bm{x}_{1})\cdots\varphi(\bm{x}_{N})\varphi^{\dagger}(\bm{x}^{\prime}_{1})\cdots\varphi^{\dagger}(\bm{x}^{\prime}_{N})\mathcal{V}_{0}^{\dagger}(\varphi)\mathcal{V}_{0}(\varphi)$
@@
$~{}\mathcal{V}_{0}(\varphi)\mathcal{V}^{\dagger}_{0}(\widetilde{\varphi})+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}~{}\frac{1}{N!}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})\widetilde{\varphi}(\bm{x}_{1})\cdots\widetilde{\varphi}(\bm{x}_{N})\mathcal{V}_{0}(\varphi)\mathcal{V}^{\dagger}_{0}(\widetilde{\varphi})$
@@
$\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)=\int\mathscr{D}\varphi~{}\frac{1}{\sqrt{N!}}\varphi(\bm{x}_{1})\cdots\varphi(\bm{x}_{N})~{}\mathcal{V}_{0}^{\dagger}(\varphi)\varPsi(\varphi,t).$ (156)
@@
$|\varPsi(t)\rangle$ $=\psi^{(0)}(t)|0\rangle+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\widehat{\varphi}^{\dagger}(\bm{x}_{1})\cdots\widehat{\varphi}^{\dagger}(\bm{x}_{N})|0\rangle$
@@
$~{}+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\int\mathscr{D}\varphi|\varphi\rangle\langle\varphi|\widehat{\varphi}^{\dagger}(\bm{x}_{1})\cdots\widehat{\varphi}^{\dagger}(\bm{x}_{N})|0\rangle$
$=\psi^{(0)}(t)|0\rangle+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N})|0\rangle$
@@
${\varPsi}(\{\varphi\},t)\equiv\psi^{(0)}(t)+\sum_{N=1}^{\infty}\int d^{3}x_{1}\cdots d^{3}x_{N}\psi(\bm{x}_{1},\cdots,\bm{x}_{N},t)\frac{1}{\sqrt{N!}}\varphi^{\dagger}(\bm{x}_{1})\cdots\varphi^{\dagger}(\bm{x}_{N}).$ (158)